### v1

Abstract
This paper describes the 7th Affective Behavior Analysis in-the-wild (ABAW) Competition, which is part of the respective Workshop held in conjunction with ECCV 2024. The 7th ABAW Competition addresses novel challenges in understanding human expressions and behaviors, crucial for the development of human-centered technologies. The Competition comprises of two sub-challenges: i) Multi-Task Learning (the goal is to learn at the same time, in a multi-task learning setting, to estimate two continuous affect dimensions, valence and arousal, to recognise between the mutually exclusive classes of the 7 basic expressions and ’other’), and to detect 12 Action Units); and ii) Compound Expression Recognition (the target is to recognise between the 7 mutually exclusive compound expression classes). s-Aff-Wild2, which is a static version of the A/V Aff-Wild2 database and contains annotations for valence-arousal, expressions and Action Units, is utilized for the purposes of the Multi-Task Learning Challenge; a part of C-EXPR-DB, which is an A/V in-the-wild database with compound expression annotations, is utilized for the purposes of the Compound Expression Recognition Challenge. In this paper, we introduce the two challenges, detailing their datasets and the protocols followed for each. We also outline the evaluation metrics, and highlight the baseline systems and their results. Additional information about the competition can be found at https://affective-behavior-analysis-in-the-wild.github.io/7th.
Keywords:
ABAW and affective behavior analysis in-the-wild and multi-task learning and compound expression recognition and Aff-Wild2 and s-Aff-Wild2 and C-EXPR-DB and valence-arousal estimation and basic-compound expression recognition and action unit detection
1 Introduction
Facial behavior analysis, bridging computer vision, physiology, and psychology, has broad applications including medicine, e-learning, marketing, entertainment and law. Recent advancements in large-scale datasets and deep learning have propelled the field forward significantly.
The 7th Affective Behavior Analysis in-the-wild (ABAW) Workshop and Competition continues to promote interdisciplinary collaboration by uniting experts from academia, industry, and government. Held alongside ECCV 2024, this workshop focuses on analyzing affective behavior in real-world contexts. This research is pivotal for advancing human-centered technologies such as human-computer interaction (HCI) systems and intelligent digital assistants. By improving our understanding of human emotions and behaviors, machines can more effectively engage with users across diverse contexts, including variations in age, gender, and social background, thereby fostering trust and improving real-life interactions.
The ABAW Competition is a crucial component of the Workshop and consists of two distinct challenges.
The first challenge focuses on Multi-Task Learning (MTL). This challenge aims to simultaneously address three tasks: (i) estimating two continuous affect dimensions (valence and arousal), (ii) recognizing eight distinct classes of facial expressions, and (iii) detecting the activation of 12 Action Units (AUs). Valence represents the positivity or negativity of an affective state on a continuous scale ranging from -1 to 1, while arousal measures the intensity of the affective state, from active to passive, also on a scale from -1 to 1. The eight expression classes include the six basic expressions (anger, disgust, fear, happiness, sadness, and surprise), the neutral state, and an ’other’ category for affective states that do not fit into the other seven classes. Action Units correspond to specific facial muscle movements, with the selected AUs for this challenge being: AU1, AU2, AU4, AU6, AU7, AU10, AU12, AU15, AU23, AU24, AU25, and AU26.
For this Challenge, a static version of the Aff-Wild2 database [118, 44, 58, 46, 61, 60, 48, 37, 47, 52, 54, 52, 38, 53, 49, 43, 40, 59, 42], named s-Aff-Wild2, is used. It consists of around  [^2] K images that contain annotations in terms of the previously mentioned tasks.
The participants are allowed to use the provided s-Aff-Wild2 database and/or any publicly available or private databases. The participants are not allowed to use the (A/V) Aff-Wild2 database (images and annotations). Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Recognition, AU detection, Face Recognition). Any methodological solution will be accepted for this Challenge.
The second Challenge targets Compound Expression (CE) Recognition. This task involves identifying seven distinct compound expressions in each frame of the provided dataset. The compound expressions to be recognized are: Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful, and Sadly Angry.
For this Challenge, a subset of the C-EXPR-DB database [38] is utilized, comprising 56 videos. C-EXPR-DB is an audiovisual (A/V) in-the-wild database containing a total of 400 videos and approximately 200,000 frames. The subset provided for this Challenge includes 56 videos with roughly 26,500 frames, which are unannotated.
Teams are allowed to use any -publicly or not- available pre-trained model and any -publicly or not- available database (that contains any annotations, e.g. VA, basic or compound expressions, AUs).
The seventh iteration of the ABAW Competition will take place as part of the ABAW Workshop, held in conjunction with ECCV 2024. This competition continues the successful legacy of previous ABAW Competitions, which have been held alongside IEEE CVPR 2024, 2023, 2022 and 2017, ECCV 2022, ICCV 2021 and IEEE FG 2020. These events have consistently attracted numerous teams from academia and industry worldwide [127, 121, 106, 92, 98, 105, 111, 130, 45, 129, 127, 92, 109, 114, 20, 77, 76, 130, 94, 36, 11, 67, 126, 14, 8, 6, 12, 123, 97, 102, 120, 108, 31, 1, 81, 64, 17, 16, 112, 71, 18, 72, 82, 29, 24, 10, 88, 73, 119, 79, 90, 33, 86, 125, 113, 35, 83, 110, 27, 104, 25, 96, 103, 30, 9, 22, 95, 7, 19, 68, 101, 78, 91, 69, 122, 65, 28, 66, 49, 75, 107, 99, 34, 93, 117, 15, 128, 84, 124, 128, 115, 70, 80, 107, 116, 87, 100, 23, 55, 45, 41, 26, 85, 39, 2, 21, 89, 4, 32, 5, 74, 3, 50, 63, 56, 57, 62, 51].
2 Competition Corpora
Below, we provide an overview of the databases used in each Challenge. For detailed information, we recommend referring to the original documentation. Additionally, we describe the pre-processing steps taken for the Challenges, which involve cropping and aligning all image frames. These steps were utilized in our baseline experiments.
2.1 Multi-Task Learning Challenge
A static version of the Aff-Wild2 database, termed s-Aff-Wild2, was created by selecting specific frames from the original database. This Challenge’s database includes 221,928 images annotated for: (i) valence and arousal; (ii) eight expressions (six basic expressions: anger, disgust, fear, happiness, sadness, and surprise; the neutral state; and an ’other’ category for all other affective states); and (iii) twelve Action Units (AUs).
Figure 1 illustrates the 2D Valence-Arousal histogram of annotations in s-Aff-Wild2. Table 1 displays the distribution of the eight expression annotations in s-Aff-Wild2. Table 2 lists the twelve annotated AUs, their corresponding actions, and their distribution in s-Aff-Wild2.
Figure 1: Multi-Task-Learning Challenge: 2D Valence-Arousal Histogram of Annotations in s-Aff-Wild2
Table 1: Multi-Task-Learning Challenge: Number of Annotated Images for each of the 8 Expressions
Expressions No of Images
Neutral 37,073
Anger 8,094
Disgust 5,922
Fear 6,899
Happiness 32,397
Sadness 13,447
Surprise 9,873
Other 39,701
Table 2: Multi-Task-Learning Challenge: Distribution of AU Annotations in Aff-Wild2
Action Unit # Action
Total Number
of Activated AUs
AU 1 inner brow raiser 29,995
AU 2 outer brow raiser 14,183
AU 4 brow lowerer 31,926
AU 6 cheek raiser 49,413
AU 7 lid tightener 72,806
AU 10 upper lip raiser 68,090
AU 12 lip corner puller 47,820
AU 15 lip corner depressor 5,105
AU 23 lip tightener 6,538
AU 24 lip pressor 8,052
AU 25 lips part 122,518
AU 26 jaw drop 19,439
The s-Aff-Wild2 dataset is divided into training, validation, and test sets. Initially, the training and validation sets, along with their annotations, are made available to participants for developing and testing their methodologies. The test set, however, is provided without annotations at a later stage.
Participants receive two versions of s-Aff-Wild2: cropped one and cropped-aligned one. All images/frames are processed using RetinaFace [13] to extract face bounding boxes and five facial landmarks. The images are cropped based on these bounding boxes. All cropped images have dimensions of $112\times 112\times 3$. The cropped-aligned version is obtained by applying similarity transformation using the five facial landmarks (two eyes, nose, and two mouth corners). Both versions are provided to participants, and the cropped-aligned version was used in our baseline experiments (described in Section 4).
2.2 Compound Expression Recognition Challenge
This Challenge utilizes a subset of the C-EXPR-DB database, consisting of 56 videos. The full C-EXPR-DB is an audiovisual in-the-wild dataset with 400 videos totaling approximately 200,000 frames, each annotated with 12 compound expressions. For this Challenge, the focus will be on recognizing the following seven compound expressions: Fearfully Surprised, Happily Surprised, Sadly Surprised, Disgustedly Surprised, Angrily Surprised, Sadly Fearful, and Sadly Angry.
Participants receive an unannotated subset of C-EXPR-DB (56 videos with about 26,500 frames) and must develop methodologies to recognize the seven specified compound expressions on a per-frame basis. Approaches may include supervised or self-supervised learning, domain adaptation, and zero- or few-shot learning techniques.
3 Evaluation Metrics for each Challenge
In this subsection, we detail the metrics used to evaluate the performance of methodologies developed by participating teams in each challenge.
3.1 Multi-Task Learning Challenge
The performance assessment for this Challenge is a composite measure that includes the following:
  • Valence and Arousal Estimation: Evaluated using the Concordance Correlation Coefficient (CCC).
  • Expression Recognition: Evaluated using the macro F1 Score across eight expression categories.
  • Action Unit Detection: Evaluated using the binary F1 Score for each of the twelve AUs.
Concordance Correlation Coefficient (CCC): CCC measures the agreement between two variables and ranges from -1 to 1, with higher values indicating better agreement. It is calculated as follows:
$\rho_{c}=\frac{2s_{xy}}{s_{x}^{2}+s_{y}^{2}+(\bar{x}-\bar{y})^{2}},$ (1)
where $s_{x}$ and $s_{y}$ are the variances of all video valence/arousal annotations and predicted values, respectively, $\bar{x}$ and $\bar{y}$ are their corresponding mean values and $s_{xy}$ is the corresponding covariance value.
F1 Score: F1 Score is the harmonic mean of precision and recall, providing a balance between the two. It ranges from 0 to 1, with higher values indicating better performance. It is defined as follows:
$F_{1}=\frac{2\times precision\times recall}{precision+recall}$ (2)
Overall Evaluation Criterion: The total performance measure for the Multi-Task Learning Challenge is the sum of the average CCC for valence and arousal, the macro F1 Score for expressions, and the binary F1 Score for AUs. The formula is as follows:
$\mathcal{P}_{MTL}$ $=\mathcal{P}_{VA}+\mathcal{P}_{EXPR}+\mathcal{P}_{AU}$
$=\frac{\rho_{a}+\rho_{v}}{2}+\frac{\sum_{expr}F_{1}^{expr}}{8}+\frac{\sum_{au}F_{1}^{au}}{12}$ (3)
3.2 Compound Expression Recognition Challenge
The evaluation metric for this challenge is based on the average F1 Score across the seven compound expressions. Thus, the performance criterion for the Compound Expression Recognition Challenge is defined as follows:
$\mathcal{P}_{CE}=\frac{\sum_{expr}F_{1}^{expr}}{7}$ (4)
This formula calculates the mean F1 Score for all seven compound expressions, ensuring a balanced evaluation of performance across the different expression categories.
4 Baseline Network and Performance
The baseline system is built entirely using open-source machine learning libraries to guarantee the reproducibility of results. Implemented in TensorFlow, the training process took approximately five hours on a Titan X GPU, utilizing a learning rate of $10^{-4}$ and a batch size of 128.
4.1 Multi-Task Learning Challenge
The baseline model utilized is a VGG16 network with its convolutional layers fixed, training only the three fully connected layers. This model was pre-trained on the VGGFACE dataset. The output layer includes 22 units: 2 linear units for valence and arousal predictions, 8 softmax units for expression predictions, and 12 sigmoid units for action unit predictions. During training with the cropped-aligned version of the s-Aff-Wild2 database, the MixAugment technique [85] was employed for data augmentation. Additionally, all image pixel intensity values were normalized to the range $[-1,1]$.
Table 3 presents the evaluation results on the s-Aff-Wild2 validation set, showing the combined score of the average CCC for valence and arousal, the average F1 score for the expression categories, and the average F1 score for the AUs.
Table 3: Multi-Task Learning Challenge results of the baseline model on the validation set of s-Aff-Wild2; overall metric is in %
Teams
Overall Metric
Github
VGGFACE with MixAugment [85] 32.0 -
VGGFACE without MixAugment [85] 30.0 -
5 Conclusion
In this paper we have presented the seventh Affective Behavior Analysis in-the-wild Competition (ABAW) held in conjunction with ECCV 2024. This Competition follows its predecessors from CVPR 2024, 2023, 2022 & 2017, ECCV 2022, ICCV 2021 and IEEE FG 2020. It features two challenges: Multi-Task Learning (MTL) to concurrently address Valence-Arousal Estimation, Expression Recognition, and Action Unit Detection; and Compound Expression Recognition, focusing on recognizing seven compound expressions. The databases utilized for this Competition are s-Aff-Wild2 and a subset of C-EXPR-DB.